OpenStudy (anonymous):

What is the pH of a solution whose [H+] is 0.0007 M?

OpenStudy (anonymous):

A particular solution has a hydrogen ion concentration of 0.008 M. What is the pH of this solution?

OpenStudy (frostbite):

Use the definition of pH \[\Large pH=-\log[H ^{+}]\]

OpenStudy (anonymous):

Frostbite can you do 1 to show me?

OpenStudy (anonymous):

Cuz I have no clue

OpenStudy (frostbite):

Okay. Lets assume we got a concentration of protons on \(1 \times 10^{-7}\). Using the definition of pH we get: \[\Large pH=-\log[H^{+}]=-\log( 1 \times 10^{-7})=-(-7)=7\]

OpenStudy (frostbite):

Try put in your own number and evaluate.

OpenStudy (frostbite):

I suggest you try calculate it your self, but answer is found below in the link: http://bit.ly/1ijNnWb